Motorboat Travel. A motorboat traveled for 2 hrs with an 8-km/hr current. The return trip against the same current took 3 hrs. Find the speed of the motorboat in still water.
I have so far but don't know if I am even close
2x + 16 = 3x + -24
2x = 3x + -24 - 16
2x = 3x + -40
-3x + 2x = -40
-x = -40
x = 40
but I not think that is all I am suppose to do.
Thank You Tammy

Answer by stanbon(29603)

Motorboat Travel. A motorboat traveled for 2 hrs with an 8-km/hr current. The return trip against the same current took 3 hrs. Find the speed of the motorboat in still water.
----------------------------
Let the speed of the boat in still water be "b".
---------------
Downstream DATA:
Time = 2 hrs ; Rate = b+8 kph ; Distance = rt = 2(b+8) km
-----------
Upstream DATA:
Time = 3 hrs ; Rate = b-8 kph ; Distance = 3(b-8)
-------------
EQUATION:
distace down = disance up
2(b+8) = 3(b-8)
2b+16 = 3b-24
b = 40 kph (speed of the boat in still water)
=================
Your answer is good if you know what you are doing.
===================
Cheers,
Stan H.